# NormXLogit: The Head-on-Top Never Lies

### v1

@@ Abstract
The Transformer architecture has emerged as the dominant choice for building large language models (LLMs). However, with new LLMs emerging on a frequent basis, it is important to consider the potential value of architecture-agnostic approaches that can provide interpretability across a variety of architectures. Despite recent successes in the interpretability of LLMs, many existing approaches rely on complex methods that are often tied to a specific model design and come with a significant computational cost. To address these limitations, we propose a novel technique, called NormXLogit, for assessing the significance of individual input tokens. This method operates based on the input and output representations associated with each token. First, we demonstrate that during the pre-training of LLMs, the norms of word embeddings capture the importance of input tokens. Second, we reveal a significant relationship between a token’s importance and the extent to which its representation can resemble the model’s final prediction. Through extensive analysis, we show that our approach consistently outperforms existing gradient-based methods in terms of faithfulness. Additionally, our method achieves better performance in layer-wise explanations compared to the most prominent architecture-specific methods.
@@ 1 Introduction
Transformer-based models have gained widespread adoption across various natural language processing (NLP) tasks, demonstrating their versatility. However, the underlying mechanisms of these models are not quite understood. This means when the model fails and generates inaccurate, toxic, or harmful content, we are unable to diagnose the source and improve the model’s behavior. Consequently, a multitude of endeavors in recent years aimed at enhancing the interpretability of these models (Kobayashi et al. 2021; Modarressi et al. 2023; Mohebbi et al. 2023).
Architecture-agnostic methods like perturbation-based and gradient-based techniques are commonly used to identify important input tokens influencing a model’s predictions. In perturbation-based methods, one can simply erase the input tokens one at a time to observe their impact on the model’s output (Li, Monroe, and Jurafsky 2017). However, a common limitation of these methods is that they can inadvertently create out-of-distribution and nonsensical inputs, which can lead to misleading results. In another direction, the main idea behind the gradient-based methods is to compute the derivative of the output with respect to input tokens to find their importance (Li et al. 2016). While gradient-based methods have been shown to be more faithful than perturbation-based methods, Wang et al. (2020) has demonstrated that they are easily manipulable and may not provide reliable interpretations.
@@
By leveraging the internal components of the target model, a new class of architecture-specific approaches, termed vector-based methods, has been developed to decompose each token into its constituent representations (Kobayashi et al. 2021; Ferrando, Gállego, and Costa-jussà 2022). Most of these approaches offer per-layer explanations, which are subsequently aggregated using the rollout technique (Abnar and Zuidema 2020) to achieve a global interpretation that integrates all layers of the model. However, this method of aggregation can result in inaccurate outcomes due to the vanishing attribution problem (Mehri et al. 2024). Despite recent advancements in faithfulness, a key drawback of vector-based methods is their architecture-specific design, which limits their adaptability to the rapidly evolving landscape of LLMs. Furthermore, almost all of these methods ignore the head-on-top111By ‘head-on-top’, we are referring to the classification or regression head used on top of the pre-trained models. which is crucial to produce task-dependent explanations.
While these methods have made significant strides, additional challenges have emerged. Improvements to these techniques often come with considerable computational costs. In perturbation-based methods, it is costly to search for appropriate combinations of tokens to intervene, as this requires multiple forward passes and adjustments to the input. Similarly, in advanced gradient-based methods like Integrated Gradients (Sundararajan, Taly, and Yan 2017) and SmoothGrad (Smilkov et al. 2017), attributions are computed by iteratively applying backpropagation. This approach can incur substantial computational overhead due to the repeated forward and backward passes required.
@@
In response to these limitations, this paper presents NormXLogit, a simple yet powerful, architecture-agnostic approach that outperforms many of the sophisticated techniques. NormXLogit does not depend on any specific architecture and could be easily applied to any NLP task. This method leverages the rich semantic and syntactic information encoded in the final layer representations of the model, in conjunction with the norm of input embeddings, resembling an end-to-end interpretation that eliminates the need for any aggregation method, such as rollout. Motivated by DecompX (Modarressi et al. 2023), we incorporate the head-on-top of pre-trained models into our analysis to obtain importance scores with respect to the task. We posit that the head-on-top of the pre-trained language models serves as a task-specific interpreter of token attribution222We use ‘attribution’ and ‘importance’ interchangeably., tailoring its attention to the nuances of the task at hand. This helps NormXLogit to generate per-label attributions with positive and negative impacts of each input token (Figure 1(a)). In addition, our method eliminates the need for any backward passes, utilizing the pre-trained model only once to generate the contextualized representation of each token, which significantly reduces the computational cost compared to other methods.
NormXLogit is characterized by combining the informativeness of input embeddings with the interpretation provided by the head-on-top. Based on comprehensive evaluation across various tasks and models, we demonstrate that in numerous instances, the faithfulness of this method surpasses that of widely recognized gradient-based approaches. Moreover, in regression setups where a classification head is absent, our method demonstrates superior performance compared to architecture-specific methods. Additionally, to evaluate layer-wise explanations, we conduct experiments on various linguistic phenomena to assess the model’s ability to differentiate between grammatically correct and incorrect sentences. Our results indicate that, in most scenarios, our proposed method performs better than or is competitive with other approaches.
@@ 2 Related Work
In recent years, vector-based analysis of Transformers has emerged as an architecture-specific approach. The motivation behind these methods stems from the observation that attention weights can be misleading for interpretability purposes (Jain and Wallace 2019; Serrano and Smith 2019). Kobayashi et al. (2021) extended the Transformer’s components for analysis by capturing the whole attention block, rather than relying solely on attention weights. In order to aggregate the per-layer attributions of previous methods to obtain a global attribution, Abnar and Zuidema (2020) proposed two different approaches to quantify the flow of information through self-attention; attention rollout and attention flow.
GlobEnc (Modarressi et al. 2022) and ALTI (Ferrando, Gállego, and Costa-jussà 2022) tried to improve previous work by further decomposing other components in the Transformer layers and aggregating them using attention rollout. ALTI questioned the way Kobayashi et al. (2021) interprets the contribution of each decomposed vector using its norm. They suggested that Manhattan distance can yield better results due to the properties of the representation space. At the local level, GlobEnc expands the scope of components beyond the attention block. They ignored the impact of the feed-froward network on the contribution of each token and tried to approximate it by adding the second residual connection and layer normalization into the analysis. In order to capture the influence of the feed-forward network, Modarressi et al. (2023) proposed DecompX. DecompX solved the problems of using rollout by propagating decomposed vectors through different layers of the Transformer. DecompX approximated and decomposed the activation function which led to a better attribution score according to their experiments. They also incorporated the classification head into the analysis which caused per-label explanations.
Perturbation-based methods typically involve perturbing or erasing parts of the input to investigate the causal relationship between input features and the model’s final prediction. This category of interpretability methods includes well-known approaches such as SHAP (Lundberg and Lee 2017) and LIME (Ribeiro, Singh, and Guestrin 2016). Recently, Mohebbi et al. (2023) proposed Value Zeroing which is based on the Explaining-by-Removing intuition (Covert, Lundberg, and Lee 2022), in order to quantify the context mixing. Their approach, in contrast to other perturbation-based methods, does not remove the input token representations. Value Zeroing instead suggests zeroing the value vector of each token to measure its contribution.
Gradient-based methods involve analyzing the gradients of the model’s output with respect to the input features to understand their impact on the model’s decision-making process. Methods such as Gradient Norm (Simonyan, Vedaldi, and Zisserman 2014), GradientInput (Kindermans et al. 2016), and Integrated Gradients (Sundararajan, Taly, and Yan 2017) are the most prominent ones in this category. However, the trustworthiness of gradient-based models has indeed been a subject of scrutiny and questioning (Adebayo et al. 2020).
@@ 3.1 Background: Transformer Architecture
Transformers are built upon an encoder-decoder structure, where the encoder and decoder components consist of multiple identical layers stacked on top of each other. In this section, we offer a concise overview of the operations carried out within an encoder layer.
An encoder layer of the Transformer architecture incorporates two sub-layers. A multi-head self-attention block (MHA) and a position-wise fully connected feed-forward network (FFN), each followed by a residual connection (RES) and layer normalization (LN). The output of each sub-layer is:
where , is the -th input representation in layer , and Sublayer() shows the functionality applied by each one of the sublayers to .
MHA.
This component is responsible for creating contextualized representations for the input elements. The output representation of MHA for token and head in -th layer is then calculated as the weighted sum of transformed input representations:
where represents the attention weight of token with respect to token in the -th head of the MHA of the -th layer. Then, in order to aggregate the outputs of all heads:
where is the number of attention heads, is the weight matrix, and is the bias vector of the final projection for the -th layer.
RES and LN.
The RES takes the input and output of the MHA and adds them together which is followed by the LN:
FFN.
The output of the first LN is passed through the FFN, which applies two linear transformations to it with a ReLU activation function in between. The input and output dimension of FFN is equal to . Then similarly, the output of the FFN passes through another LN and undergoes a RES:
The and are the parameters corresponding to the first and second linear transformations of the -th layer.
@@ 3.2 Norm of Word Embedding
Oyama, Yokoi, and Shimodaira (2023) demonstrated that the norm of input embeddings encodes information gain. They showed that tokens with higher norm carry more information, effectively capturing the least frequent words in the text. Additionally, based on the Eq. 2, the MHA could be interpreted as the weighted sum of transformed vectors. In other words, the final representation of each token is built by mixing the representations of all tokens in the input sequence. Consequently, tokens with higher norms are expected to contribute more to the final representation of the target token. Higher contribution suggests greater importance, allowing us to utilize the norm of word embeddings to identify crucial tokens influencing the model’s decision.
@@ 3.3 LogAt: Logit Attribution
The tasks in the domain of NLP can be broadly divided into two main categories: classification tasks and regression tasks. For both of these setups, we utilize a special token (often known as [CLS]333The name of this special token may vary depending on the model. Also, in auto-regressive models, the last token in the input is typically used for classification.), which is embedded in almost all pre-trained models. This token serves as a single vector representing the entire input sequence, which is then fed into head-on-top, an FFN placed on top of the pre-trained model to produce the output prediction.
@@ 4.1 Experimental Setup
To analyze the faithfulness of NormXLogit we conduct our first experiment on classic classification and regression tasks.
@@
In the classification setup, we will use SST-2 (Socher et al. 2013) for sentiment analysis and MultiNLI (Williams, Nangia, and Bowman 2018) for recognizing textual entailment. SST-2 contains sentences with negative and positive labels extracted from movie reviews, while MultiNLI includes sentence pairs labeled as entailment, contradiction, and neutral. Additionally, we employ STS-B (Cer et al. 2017) to evaluate semantic textual similarity as a regression task. This dataset provides a benchmark for measuring the similarity between sentence pairs, with annotations ranging from 0 (no similarity) to 5 (semantic equivalence).
@@
Our target models in this section involve three prominent models: Llama 2 (Touvron et al. 2023), DeBERTa (He, Gao, and Chen 2023), and BERT (Devlin et al. 2019).444Specifically, we employ the 7 billion parameter variant of Llama 2, the uncased BERTbase model, and DeBERTaV3base, all of which are obtained from HuggingFace’s Transformers library (Wolf et al. 2020). We use the fine-tuned version of each model for the corresponding task. To fine-tune Llama 2 and perform inference, we employ the LoRA (Hu et al. 2021) technique with a rank of 4 from the PEFT library555https://github.com/huggingface/peft.
| SST-2 (AOPC) | MNLI (AOPC) | STS-B (Acc) | |||||||
| Random Baseline | 0.256 | 0.266 | 0.245 | 0.421 | 0.445 | 0.361 | 0.283 | 0.430 | 0.457 |
| Gradient Norm | 0.216 | 0.320 | 0.331 | 0.419 | 0.535 | 0.460 | 0.351 | 0.338 | 0.374 |
| Gradient×Input | 0.236 | 0.345 | 0.339 | 0.442 | 0.565 | 0.456 | 0.255 | 0.214 | 0.358 |
| Integrated Gradients | 0.220 | 0.346 | 0.367 | 0.448 | 0.571 | 0.466 | 0.237 | 0.227 | 0.370 |
| DecompX | N/A | N/A | 0.574 | N/A | N/A | 0.585 | N/A | N/A | 0.336 |
| norm | 0.299 | 0.360 | 0.311 | 0.420 | 0.473 | 0.393 | 0.251 | 0.199 | 0.321 |
| LogAt | 0.341 | 0.377 | 0.364 | 0.518 | 0.548 | 0.566 | 0.167 | 0.423 | 0.313 |
| NormXLogit | 0.341 | 0.386 | 0.423 | 0.519 | 0.566 | 0.556 | 0.233 | 0.320 | 0.281 |
@@
To analyze the performance of our proposed method, we compare NormXLogit with three well-known gradient-based input attribution methods: Gradient Norm, GradientInput, and Integrated Gradients. For all of these methods, we use the norm as the aggregation approach. To account for vector-based approaches, we adopt DecompX as the current state-of-the-art method among them. However, this family of methods is primarily developed for BERT-like architectures and may not be applicable to all models. In addition, we consider a random baseline where tokens are ranked randomly from most important to least important.
@@
To assess the faithfulness of the aforementioned methods, we utilize two metrics: AOPC (Samek et al. 2015) for classification tasks and Accuracy for regression setups.
@@
This metric involves perturbing of the most important tokens in the input sequence and observing the resulting changes in the model’s predictions. For the masked language modeling objectives, masking is used for token perturbations, while for auto-regressive models, deletions are employed due to the absence of a [MASK] token. For a given input sequence , the perturbed sequence is generated by applying the perturbation on of the most important tokens. Then for all of the instances in the dataset, the average AOPC is defined as:
where is the number of instances, and is the model’s output probability for label . A higher AOPC indicates that the model exhibits a larger drop in probability for the predicted class, reflecting greater sensitivity to perturbated tokens.
@@
This metric operates by observing the accuracy drop when perturbing different proportions of the most important tokens in the input. For regression tasks, we utilized the Pearson correlation coefficient as the accuracy metric. For the Accuracy, lower values indicate better performance.
@@ 4.2 Results
Table 1 presents the average AOPC and Accuracy across different ratios of perturbation, evaluated on various models and datasets666The corresponding diagrams are also provided in the Appendix.. In SST-2, NormXLogit consistently outperforms architecture-agnostic methods. However, DecompX, which is specific to the BERT architecture, results in a higher drop in AOPC. In the MultiNLI dataset, NormXLogit performs better than gradient-based approaches in Llama 2 and BERT, though Integrated Gradients show a slight edge in the DeBERTa model. In the BERT model, similar to SST-2, DecompX performs better but the difference is notably smaller compared to the SST-2 dataset.
In the regression setup, surprisingly, the norm outperforms other methods and also helps GradientInput achieve better results. In the BERT model, the absence of a classification head diminishes DecompX’s effectiveness, resulting in performance worse than that of the input embeddings’ norms. In Llama 2, NormXLogit slightly surpasses all gradient-based methods, largely due to the strong performance of LogAt.

### v2

@@ Abstract
With new large language models (LLMs) emerging frequently, it is important to consider the potential value of model-agnostic approaches that can provide interpretability across a variety of architectures. While recent advances in LLM interpretability show promise, many rely on complex, model-specific methods with high computational costs. To address these limitations, we propose NormXLogit, a novel technique for assessing the significance of individual input tokens. This method operates based on the input and output representations associated with each token. First, we demonstrate that during the pre-training of LLMs, the norms of word embeddings effectively capture token importance. Second, we reveal a significant relationship between a token’s importance and the extent to which its representation can resemble the model’s final prediction. Extensive analyses reveal that our approach outperforms existing gradient-based methods in terms of faithfulness and offers competitive performance in layer-wise explanations compared to leading architecture-specific techniques.
NormXLogit: The Head-on-Top Never Lies
Sina Abbasi1 Mohammad Reza Modarres1 Mohammad Taher Pilehvar2 1 Tehran Institute for Advanced Studies, Khatam University, Iran 2 Cardiff University, United Kingdom s.abbasi401@khatam.ac.ir m.modares401@khatam.ac.ir mp792@cam.ac.uk
@@ 1 Introduction
Transformer-based models have gained widespread adoption across various natural language processing (NLP) tasks, demonstrating their versatility. However, the underlying mechanisms of these models are not quite understood. This means when the model fails and generates inaccurate or harmful content, we are unable to diagnose the source and improve the model’s behavior. Consequently, a multitude of endeavors in recent years aimed at enhancing the interpretability of these models.
Architecture-agnostic methods, such as perturbation-based and gradient-based techniques, are widely used to identify influential input tokens that impact a model’s predictions. However, these approaches face significant challenges, including the generation of out-of-distribution inputs and vulnerability to adversarial exploitation (Wang et al., 2020). Additionally, they often come with high computational costs, limiting their scalability. For instance, advanced gradient-based methods like Integrated Gradients (Sundararajan et al., 2017) involve repeated forward and backward passes, leading to substantial overhead.
@@
By leveraging the internal components of the target model, a new class of architecture-specific approaches, termed vector-based methods, has been developed (Kobayashi et al., 2021; Ferrando et al., 2022). Most of these approaches offer per-layer explanations, which are subsequently aggregated using the rollout technique (Abnar and Zuidema, 2020) to achieve a global interpretation that integrates all layers of the model. However, rollout can result in inaccurate outcomes due to the vanishing attribution problem (Mehri et al., 2024). Despite recent advances in faithfulness, a key drawback of vector-based methods is their architecture-specific design, which limits their adaptability to the rapidly evolving landscape of LLMs. Furthermore, almost all of these methods ignore the head-on-top111By ‘head-on-top’, we are referring to the classification or regression head used on top of the pre-trained models. which is crucial to produce task-dependent explanations.
@@
In response to these limitations, this paper presents NormXLogit, a simple yet powerful, computationally efficient, and architecture-agnostic222By using the term ”architecture-agnostic,” we aim to differentiate our approach from vector-based methods, ensuring generalization across different Transformer-based models. approach that outperforms many sophisticated techniques and can be easily applied to any NLP task. NormXLogit eliminates the need for any backward passes, utilizing the pre-trained model only once. It leverages the informativeness of the norm of input embeddings, in conjunction with the rich semantic and syntactic information encoded in the model’s final-layer representations. The latter enables the incorporation of all internal components, thereby removing the need for any aggregation method, such as rollout. We incorporate the head-on-top of pre-trained models into our analysis, which serves as a task-specific interpreter of token attributions333We use ‘attribution’ and ‘importance’ interchangeably., tailoring its attention to the nuances of the task. This helps NormXLogit to generate per-label attributions with positive and negative impacts of each input token (cf. Figure 1(a)).
Based on comprehensive experiments across multiple tasks and models, we show that in numerous instances, the faithfulness of NormXLogit surpasses that of widely recognized gradient-based approaches. In the regression setup, it demonstrates superior performance compared to a state-of-the-art architecture-specific baseline. Furthermore, through a targeted experiment on the BLiMP corpus (Warstadt et al., 2020), we evaluate the alignment between NormXLogit’s attributions and known linguistic evidence across diverse grammatical phenomena. Our results demonstrate that NormXLogit consistently yields more faithful and plausible explanations than existing approaches, highlighting its effectiveness as a general-purpose token attribution method.
@@ 2 Related Work
In recent years, vector-based analyses have emerged as an architecture-specific approach to Transformer interpretability. These methods build on findings that attention weights can be misleading (Jain and Wallace, 2019; Serrano and Smith, 2019). Kobayashi et al. (2021) extended analysis beyond attention weights by decomposing the entire attention block. To aggregate per-layer attributions into global ones, Abnar and Zuidema (2020) proposed attention rollout to quantify the flow of information through self-attention.
GlobEnc (Modarressi et al., 2022) and ALTI (Ferrando et al., 2022) advanced previous work by further incorporating other Transformer components. ALTI challenged the use of norms for measuring contributions, proposing the Manhattan distance for improved results. GlobEnc added the second residual connection and layer normalization into the analysis but overlooked the impact of the feed-forward network. To address this, DecompX (Modarressi et al., 2023) captured the influence of the feed-forward network by approximately decomposing the activation function and propagating decomposed vectors across layers, eliminating reliance on aggregation methods like rollout.
Perturbation-based methods, such as SHAP (Lundberg and Lee, 2017) and LIME (Ribeiro et al., 2016), investigate the causal link between input features and the model’s final prediction by perturbing or erasing parts of the input. Recently, Mohebbi et al. (2023) proposed Value Zeroing which is based on the Explaining-by-Removing intuition (Covert et al., 2022), in order to quantify the context mixing. Their approach, in contrast to other perturbation-based methods, does not remove the input token representations. Value Zeroing instead suggests zeroing the value vector of each token to measure its contribution.
Gradient-based methods involve analyzing the gradients of the model’s output with respect to the input features to understand their impact on the model’s decision-making process. Methods such as Gradient Norm (Simonyan et al., 2014), GradientInput (Kindermans et al., 2016), and Integrated Gradients (Sundararajan et al., 2017) are the most prominent ones in this category.
@@ 3.1 Background: Transformer Architecture
The Transformer architecture consists of multiple identical layers, each with a multi-head self-attention (MHA) block and a position-wise fully connected feed-forward network (FFN), where both are followed by a residual connection (RES) and layer normalization (LN).
The MHA component is responsible for creating contextualized representations for the input elements. The output representation of MHA for token and head in -th layer is then calculated as the weighted sum of transformed input representations:
where represents the attention weight of token with respect to token in the -th head of the MHA of the -th layer.
@@ 3.2 Norm of Word Embedding
Oyama et al. (2023) demonstrated that the norm of input embeddings encodes information gain. They showed that tokens with higher norm carry more information, effectively capturing the least frequent words in the text. Additionally, based on the Eq. 1, the MHA could be interpreted as the weighted sum of transformed vectors. In other words, the final representation of each token is built by mixing the representations of all tokens in the input sequence. Consequently, tokens with higher norms are expected to contribute more to the final representation of the target token. Higher contribution suggests greater importance, allowing us to utilize the norm of word embeddings to identify crucial tokens influencing the model’s decision.
@@ 3.3 LogAt: Logit Attribution
The tasks in the domain of NLP can be broadly divided into two main categories: classification tasks and regression tasks. For both of these setups, we utilize a special token (often known as [CLS]444The name of this special token may vary depending on the model. Also, in auto-regressive models, the last token in the input is typically used for classification.), which is embedded in almost all pre-trained models. This token serves as a single vector representing the entire input sequence, which is then fed into head-on-top, an FFN placed on top of the pre-trained model to produce the output prediction.
@@ 4 Experiment 1: Faithfulness Analysis
To analyze the faithfulness of NormXLogit, we begin with a set of experiments on classic classification and regression tasks. The concept is straightforward: if a token is identified as highly important, altering or removing it should significantly impact the model’s output prediction. A faithful attribution method should effectively identify these critical tokens, leading to a more substantial drop in the model’s performance.
@@ 4.1 Experimental Setup
In the classification setup, we will use SST-2 (Socher et al., 2013) for sentiment analysis, MultiNLI (Williams et al., 2018) for recognizing textual entailment, and QNLI for question answering (Wang et al., 2018). SST-2 contains movie review sentences labeled as positive or negative, while MultiNLI includes sentence pairs labeled as entailment, contradiction, or neutral. QNLI consists of question–sentence pairs labeled for entailment. STS-B (Cer et al., 2017) is used for semantic similarity as a regression task, with scores from 0 (no similarity) to 5 (semantic equivalence).
@@
Our target models in this section involve three prominent models: Llama 2 (Touvron et al., 2023), DeBERTa (He et al., 2023), and BERT (Devlin et al., 2019).555We employ the 7 billion parameter variant of Llama 2, the uncased BERTbase model, and DeBERTaV3base, all from HuggingFace’s Transformers library Wolf et al. (2020). We use the fine-tuned version of each model for the corresponding task. To fine-tune Llama 2 and perform inference, we employ the LoRA (Hu et al., 2021) technique with a rank of 4 from the PEFT library666https://github.com/huggingface/peft.
@@
To analyze the performance of our proposed method, we compare NormXLogit with three well-known gradient-based input attribution methods: Gradient Norm (Grad. Norm), GradientInput (GI), and Integrated Gradients (IG), using the norm for aggregation. Notably, we focus on gradient-based methods over perturbation-based approaches due to their more faithful results. To account for vector-based approaches, we adopt DecompX as the current state-of-the-art method among them. However, this family of methods is primarily developed for BERT-like architectures and may not be applicable to all models. In addition, we consider a random baseline where tokens are ranked randomly from most important to least important.
@@
To assess the faithfulness of the aforementioned methods, we utilize two metrics: AOPC777AOPC stands for “Area Over the Perturbation Curve.” (Samek et al., 2015) for classification tasks and Accuracy for regression setups.
@@
This metric involves perturbing of the most important tokens in the input sequence and observing the resulting changes in the model’s predictions. For the masked language modeling objectives, masking is used for token perturbations, while for auto-regressive models, deletions are employed due to the absence of a [MASK] token. For a given input sequence , the perturbed sequence is generated by applying the perturbation on of the most important tokens. Then, for all of the instances in the dataset, the average AOPC is defined as:
where is the number of instances, and is the model’s output probability for label . A higher AOPC for the predicted class indicates that the model exhibits a larger drop in its probability, reflecting greater sensitivity to perturbed tokens.
@@
This metric evaluates model performance when different proportions (e.g., , , …) of the most important input tokens are perturbed. For regression tasks, we utilize the Pearson correlation coefficient as the Accuracy metric. For classification tasks, we report standard accuracy (i.e., the percentage of correctly predicted labels) on the perturbed inputs. In both cases, lower values indicate better attribution performance, as the model is more affected by key perturbations.
@@ 4.2 Results
| SST-2 (AOPC) | MNLI (AOPC) | QNLI (AOPC) | STS-B (Acc) | |||||||||
| Llama 2 | DeBERTa | BERT | Llama 2 | DeBERTa | BERT | Llama 2 | DeBERTa | BERT | Llama 2 | DeBERTa | BERT | |
| Random | 0.256 | 0.266 | 0.245 | 0.421 | 0.445 | 0.361 | 0.284 | 0.306 | 0.273 | 0.283 | 0.430 | 0.457 |
| Grad. Norm | 0.216 | 0.320 | 0.331 | 0.419 | 0.535 | 0.460 | 0.334 | 0.365 | 0.360 | 0.351 | 0.338 | 0.374 |
| GI | 0.236 | 0.345 | 0.339 | 0.442 | 0.565 | 0.456 | 0.353 | 0.382 | 0.364 | 0.255 | 0.214 | 0.358 |
| IG | 0.220 | 0.346 | 0.367 | 0.448 | 0.571 | 0.466 | 0.336 | 0.381 | 0.364 | 0.237 | 0.227 | 0.370 |
| DecompX | N/A | N/A | 0.574 | N/A | N/A | 0.585 | N/A | N/A | 0.460 | N/A | N/A | 0.336 |
| norm | 0.299 | 0.360 | 0.311 | 0.420 | 0.473 | 0.393 | 0.272 | 0.339 | 0.304 | 0.251 | 0.199 | 0.321 |
| LogAt | 0.341 | 0.377 | 0.364 | 0.518 | 0.548 | 0.566 | 0.378 | 0.435 | 0.394 | 0.167 | 0.423 | 0.313 |
| NormXLogit | 0.341 | 0.386 | 0.423 | 0.519 | 0.566 | 0.556 | 0.363 | 0.474 | 0.402 | 0.233 | 0.320 | 0.281 |
Table 1 presents the average AOPC and Accuracy across different ratios of perturbation, evaluated on various models and datasets888The corresponding diagrams are provided in Section A.2.. For classification tasks, the Accuracy metric results are consistent with AOPC scores and are provided in the appendix to avoid repetition (cf. Table 3). In SST-2 and QNLI, NormXLogit consistently outperforms architecture-agnostic methods. However, DecompX, which is specific to the BERT architecture, results in a higher drop in AOPC. In the MultiNLI dataset, NormXLogit performs better than gradient-based approaches in Llama 2 and BERT, though Integrated Gradients show a slight edge in the DeBERTa model. In the BERT model, similar to SST-2, DecompX performs better but the difference is notably smaller compared to the SST-2 dataset.
In the regression setup, the norm performs surprisingly well and also boosts the performance of GradientInput. This strong performance on STS-B can be attributed to the structure of STS-B samples: sentence pairs consist of relatively short sentences with similar openings, where key information that determines the label often resides in less frequent words toward the end. Such words tend to have higher embedding norms, making the norm approach more effective.
Furthermore, in the BERT model, the absence of a classification head diminishes DecompX’s effectiveness, resulting in performance worse than that of the input embeddings’ norms. In Llama 2, NormXLogit slightly surpasses all gradient-based methods, largely due to the strong performance of LogAt.
Computational Efficiency.
NormXLogit is not only faithful and generalizable, but also computationally lightweight. In perturbation-based methods, it is costly to search for appropriate combinations of tokens to intervene, as this requires multiple forward passes. Similarly, advanced gradient-based methods like Integrated Gradients incur substantial overhead due to repeated forward and backward passes. In contrast, NormXLogit performs attribution in a single forward pass, eliminating the need for any backpropagation. Our runtime analysis shows that for longer inputs (e.g., 360 tokens or more), NormXLogit can be up to faster and up to more memory-efficient than its counterparts (cf. Tables 4 and 5).
5 Experiment 2: Evidence Alignment
In our second experiment, we extend our evaluation beyond classification and regression by focusing on the language modeling task, using the BLiMP dataset Warstadt et al. (2020). This experiment serves four purposes: (1) to assess the faithfulness of NormXLogit in the masked language modeling setting, expanding the scope of our evaluation, (2) to examine its plausibility, i.e., the extent to which its attributions align with human-understood linguistic rationales provided by BLiMP, (3) to analyze the target model’s sensitivity to specific linguistic phenomena, and (4) to investigate how contextualization evolves throughout the model by analyzing the output representations at different layers of the Transformer (i.e., per-layer attributions).
5.1 Experimental Setup
Data.
The BLiMP dataset contains sentence pairs with minimal contrasts in syntax, morphology, or semantics. It is constructed to provide samples where the true label is uniquely determined by a single word in each sentence, providing an ideal benchmark for assessing attribution methods. This word, which serves as the decisive factor in determining grammatical acceptability, is termed the evidence. BLiMP offers a strong prior on which token is expected to drive the model’s prediction, thereby supporting a targeted assessment of attribution faithfulness. Additionally, it enables evaluation of plausibility, as the evidence in this dataset inherently aligns with human reasoning. In summary, using BLiMP we are evaluating both the faithfulness and plausibility of these methods.